### v1

# Fairness and Bias in Multimodal : A Survey

Abstract
The importance of addressing fairness and bias in artificial intelligence (AI)systems cannot be over-emphasized. Mainstream media has been awashed with news of incidents around stereotypes and bias in many of these systems in recent years. In this survey, we fill a gap with regards to the minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models along with the challenges affecting them; we identify a new category of quantifying bias (preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic; we critically discuss the various ways researchers are addressing these challenges. Our method involved two slightly different search queries on Google Scholar, which revealed that 33,400 and 538,000 links are the results for the terms "Fairness and bias in Large Multimodal Models" and "Fairness and bias in Large Language Models", respectively. We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenge of fairness and bias in multimodal AI.
1 Introduction
Fairness and bias are very important topics that cut across many domains in the society. The rapid advancements in the research and applications of artificial intelligence (AI)have made them even more compelling in recent times, such that many studies have emerged on them Frankel and Vendrow (2020); Booth et al. (2021); Adewumi et al. (2022); Teo et al. (2024). One important gap in the literature, however, is that there is minimal study or survey on "Fairness and bias in Large Multimodal Models." By multimodal AI, we mean the dataset or AI model that can take one or more modalities as input and/or another as output. This gap is evidenced by the fact that there is a dearth of work on the topic. For example, an exact-phrase search on Google Scholar returns 33,400 links compared to 538,000 links for "Fairness and bias in Large Language Models." [^1] June 10, 2024 This implies more than 16 times the result for the former. Filtering the publication year range to 2014-2024 reduces the links to 17,700 and 20,300, respectively. We intend to contribute in filling that gap in this work.
The two terms, fairness and bias, are strongly related but fairness is concerned with equality and justice while bias is concerned with systematic error, which may arise from human prejudices Booth et al. (2021); Alkhaled et al. (2023). For the purpose of this survey, fairness may be defined as equal representation with regards to a given Sensitive Attribute (SA)Hutchinson and Mitchell (2019); Frankel and Vendrow (2020). Hence, we may consider a generative artifical intelligence (GenAI)to be fair if it generates both male and female samples with equal probabilities, with regards to the SA gender Teo et al. (2024). Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one group compared to another, given the ground truth Booth et al. (2021); Scheuneman (1979). We acknowledge there are other quantitative definitions of fairness and bias, as noted by Hutchinson and Mitchell (2019) and Weidinger et al. (2022).
It appears the emergence of big data, which has brought rapid advancement in the state-of-the-art (SotA), also brought along the increase in poor quality content and prediction, such as the increased criminal prediction for Black and Latino people observed by Birhane et al. (2024a). Similar issues are observed across many domains, including healthcare, employment, forensics, criminal justice, credit scoring, and computational social science, among others Liang et al. (2021); Ferrara (2023); Landers and Behrend (2023); Han (2023). According to Wolfe et al. (2023), the model VQGAN-CLIP, similarly to Stable Diffusion, generated sexualized images for the harmless prompt "a 17 year old girl," 73% of the time. The comparison to a similar prompt with the term "girl" replaced with "boy" shows a sharp contrast.
In view of the foregoing gap and challenges, this work critically surveys the literature with the primary objective of ascertaining what the state of work is on fairness and bias in multimodal AI, thereby making the following key contributions.
  1. 1.
    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges affecting them.
  2. 2.
    We identify a new category of quantifying bias (i.e. preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic Ramesh et al. (2023); Cabello et al. (2023).
  3. 3.
    We critically discuss many important approaches to addressing the challenges of fairness and bias.
The rest of this paper is organised as follows. In the following Section (2), we discuss fairness and bias in AI, highlighting some of the theories. In Section 3, we explain the method used for this survey. In Section 4, we focus on various works along the two paradigms of LMM s and LLM s, discussing the datasets and models in the literature. In Section 5, we discuss widely on the methods to evaluate fairness and bias, the datasets for evaluation, and the debiasing strategies. We conclude the survey in Section 6 with a summary and possible future work.
2 Fairness and Bias in AI
2.1 Concepts of Fairness and Bias
A discussion about every possible type of fairness or bias is beyond the scope of this study, however, we refer readers to Mehrabi et al. (2021) for some of the discussion. In this section, we highlight some of the concepts around fairness and bias found in the Social Science literature.
2.1.1 Justice Theory
The theory is regarded as a three-part framework of distributive, interactional, and procedural justice perceptions Greenberg (1990); Landers and Behrend (2023). Distributive justice, when outcomes are expected to be distributed equally, is the overarching aspect related to AI fairness and bias, according to Landers and Behrend (2023). It is based on equality rules, need, or equity, which are influenced by social and cultural values.
2.1.2 Equity Theory
Equity theory uses a unidimensional concept of fairness instead of multidimensional, according to Leventhal (1980). It perceives justice solely on the merit principle and the final distribution of reward (or punishment), where reward is proportional to contribution Adams and Freedman (1976).
2.1.3 Objectification Theory
Just as inanimate objects have no emotions or thoughts, objectification theory establishes a view of a subject as primarily without human characteristics, especially for women and girls Fredrickson and Roberts (1997); Heflick et al. (2011); Andrighetto et al. (2019). The theory identifies sexual objectification bias, which is when the emotions or thoughts of a person are disregarded and one is treated as mere body parts for sex Fredrickson and Roberts (1997); Wolfe et al. (2023).
2.2 Consequences & Legal Perspectives
Fredrickson and Roberts (1997) confirms that objectification victimises the subject and may result in habitual body monitoring, thereby increasing mental health risks, sexual dysfunction, eating disorders, and depression. This unhealthy reality is also confirmed by Swim et al. (2001). They realized that sexist incidents occur more against women and have negative emotional consequences for them. Some of these incidents are traditional gender role stereotypes, degrading remarks, and sexual objectification, which are found in the data used for training AI models. For details about the mechanics of training language models, we refer readers to Radford et al. (2019) and Hoffmann et al. (2022). It is not surprising, therefore, that the use of these models cause the same negative effects for those affected. Hence, fairness and bias are not only ethical or moral issues but have legal implications Landers and Behrend (2023). In the United States (US), disparate treatment because of sensitive attributes is unlawful Berry (2015); Hutchinson and Mitchell (2019); Meng et al. (2022). This is also the case in some other countries Zafar et al. (2017).
3 Method
To address the gap identified earlier, our method involved the simple approach of filtered search on Google scholar with two slightly differently-worded phrases: (1) "Fairness and Bias in Large Multimodal Models" and (2) "Fairness and Bias in Large Language Models" and they returned 17,700 and 20,300 links,1 respectively, over many pages. The search was filtered to the period 2014-2024 during which deep learning made significant progress. Apparently, there are fewer scientific papers on the former. We compare with "Language Models" only because these, with the introduction of the Transformer architecture Vaswani et al. (2017), have influenced computer vision Yuan et al. (2021) and serve as core component for many multimodal AI.
We captured archived papers (e.g. arXiv) but then thoroughly reviewed pair-reviewed papers that were returned on the first 10 pages (100 links) for each search text and identified the publishers (or venues), as listed in Table 1. We note that the more relevant papers turn up on the first page of the search while the quality of results degrade as one progresses through the pages. From the Table, it can be observed that there are 36 pair-reviewed papers on multimodal compared to 39 on language, among the 200 papers. We note that Google Scholar sometimes returns the arXiv version of a paper instead of the peer-reviewed version, hence, it’s very useful to include the data from arXiv. In total, 55 papers were published under multimodal compared to 88 for language. Each paper was reviewed for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
Table 1: Distribution of scientific papers (2014-2024)
# Publisher Multimodal Language
1 IEEE 11 2
2 ELSEVIER 2 2
3 ACM 9 9
4 Springer 1 1
5 NeurIPS 4 8
6 Nature 3 1
7 MDPI - 1
8 MLR 3 5
9 PKP 2 -
10 MIT Press - 4
11 PubMed 1 -
12 Cambridge - 1
13 Patterns - 1
14 PLOS - 1
15 JMLR - 2
16 Science - 1
Sub-total 36 39
17 arXiv 19 49
Total 55 88
4 Fairness and Bias in LLM s and LMM s
In this section we review some LMM s and LLM s and the fairness and bias challenges they have. Tables 2 and 3 summarize some relevant datasets and the models, respectively. All the 25 datasets identified have their challenges and by extension the 25 AI models they are trained on. Some of these challenges include stereotypes, porn, misogyny, racial, gender, religious, cultural, age, and demographic biases. It is commonly agreed AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022).
4.1 LMM s
Table 2: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the second part of the table are usually used in downstream tasks).
# Dataset Modality Some Challenges
1 CommonCrawl Raffel et al. (2020) Text & Vision Fake news, hate speech, porn & racism Gehman et al. (2020); Luccioni and Viviano (2021)
2 LAION-400M & 5B Schuhmann et al. (2021, 2022) Text & Vision Misogyny, stereotypes & porn Birhane et al. (2021, 2024b)
3 WebImageText (WIT)Radford et al. (2021) Text & Vision Racial, gender biases Radford et al. (2021)
4 DataComp Gadre et al. (2024) Text & Vision Racial bias Gadre et al. (2024)
5 WebLI Chen et al. (2022) Text & Vision Age, racial, gender biases & stereotypes Chen et al. (2022)
6 CC3M-35L Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
7 COCO-35L Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
8 WIT Srinivasan et al. (2021) Text & Vision Cultural bias Srinivasan et al. (2021)
9 Colossal Cleaned CommonCrawl (C4) Raffel et al. (2020) Text Offensive language, racial bias Raffel et al. (2020)
10 The Pile Gao et al. (2020) Text Religious, racial, gender biases Gao et al. (2020)
11 CCAligned El-Kishky et al. (2020) Text Porn, racial bias El-Kishky et al. (2020)
12 OpenAI WebText Radford et al. (2019) Text Gender, racial biases Gehman et al. (2020)
13 OpenWebText Corpus (OWTC) Text Gender, racial biases Gehman et al. (2020)
14 ROOTS Laurençon et al. (2022) Text Cultural bias Laurençon et al. (2022)
15 VoxCeleb 1 Nagrani et al. (2020) Audio & Vision Demographic, gender biases Chung et al. (2018)
16 VoxCeleb 2 Chung et al. (2018) Audio & Vision Demographic, gender biases Chung et al. (2018)
17 First Impressions Escalante et al. (2020) Audio & Vision Racial, gender biases Yan et al. (2020)
18 XM3600 Thapliyal et al. (2022) Text & Vision Cultural bias Thapliyal et al. (2022)
19 VQA Antol et al. (2015) Text & Vision Gender bias Ruggeri and Nozza (2023)
20 VQA 2 Goyal et al. (2017) Text & Vision Gender bias Ruggeri and Nozza (2023)
21 MS COCO Lin et al. (2014) Text & Vision Gender bias Cabello et al. (2023); Zhao et al. (2017)
22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022)
23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022)
24 MAB Alkhaled et al. (2023) Text Racism, misogyny, stereotypes Alkhaled et al. (2023); Pagliai et al. (2024)
25 Twitter corpus Huang et al. (2020) Text Age, gender, racial biases Huang et al. (2020)
Wolfe et al. (2023) found evidence of sexual objectification bias in models based on Contrastive Language-Image Pretraining (CLIP). The 9 CLIP models that were investigated were trained on internet-wide web crawls. CLIP is known to be quite accurate Radford et al. (2021), however, it also appears to have scaled the biases inherent in its training data. Also, Wolfe and Caliskan (2022) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al. (2021), SLIP Mu et al. (2022), and BLIP Li et al. (2022), as measured with EATs. For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used Chicago Face Database (CFD)Ma et al. (2015). CFD is a dataset of 597 people recruited in the U.S. Similarly, Teo et al. (2024) found that Stable Diffusion exhibits gender bias on slight changes to the prompts.
Besides work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022). In the work by Fenu and Marras (2022), they perform comparative analysis on audio-visual speaker recognition systems, using fusion at the model step. They found that the highest accuracy and the lowest disparity across groups are achieved compared to unimodal systems.
In other works, Peña et al. (2023) evaluated AI-based recruitment for multimodal data but in a fictitious case study, which may limit its generalizability in real-world applications. Booth et al. (2021) performed a case study of automated video interviews and found that combining more than one modality increases bias and reduces fairness, similarly to what happens when scaling crawled data for training models. Cabello et al. (2023) investigated fairness and bias in LMM s. Other researchers investigated the impact of multimodal data/models on personlity assessment Yan et al. (2020), cyberbullying Alasadi et al. (2020), health records Meng et al. (2022) and more Birhane et al. (2021); Ferrara (2023); Zhang et al. (2022)
4.2 LLM s
Table 3: Summary of Models and Some of Their Fairness & Bias Challenges. (*modified for audio-visual)
# LMM s Modality Training Data Some Challenges
1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022)
2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023)
3 GLIDE Nichol et al. (2022) Text & Vision WIT & Conceptual Captions Gender stereotypes Nichol et al. (2022)
4 Stable Diffusion Rombach et al. (2022) Text & Vision LAION-5B Gender bias Teo et al. (2024)
5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022)
6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022)
7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022); Ruggeri and Nozza (2023)
8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language222ai.google.dev/gemma/docs/paligemma/model-card
9 PaLI-3 Chen et al. (2022) Text & Vision WebLI Age, racial, gender biases & stereotypes Chen et al. (2022)
10 Falcon 2 Almazrouei et al. (2023) Text & Vision RefinedWeb Harmful content, cultural bias Almazrouei et al. (2023)
11 BEiT Wang et al. (2023) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
12 LLaVA Liu et al. (2024b, a) Text & Vision Conceptual Captions Cultural bias Liu et al. (2024b)
13 ResNet-50* He et al. (2016) Audio & Vision ImageNet Gender bias Fenu and Marras (2022)
14 GPT4o Achiam et al. (2023) Text, Audio & Vision WebText, Github, etc Stereotypes, racial bias Aich et al. (2024)
LLM s
15 GPT3 Brown et al. (2020) Text CommonCrawl & WebText Gender, racial, religious biases Brown et al. (2020); Gehman et al. (2020)
16 PaLM Chowdhery et al. (2024) Text Wikipedia, social media, Github Occupation, gender, sexual, religious biases Chowdhery et al. (2024)
17 LaMDA Thoppilan et al. (2022) Text Social media & Wikipedia Gender bias Thoppilan et al. (2022)
18 GLaM Du et al. (2022) Text Wikipedia & social media Toxicity, gender bias Du et al. (2022)
19 GPT2 Radford et al. (2019) Text WebText Sexual, racial, gender biases Sheng et al. (2019); Gehman et al. (2020)
20 LLaMA-3 Text web text & Github Stereotypes, gender, racial, sexual, religious, biasesAich et al. (2024); Touvron et al. (2023)
21 LLaMA-2 Touvron et al. (2023) Text web text Toxicity, gender, racial, sexual, religious, biases Touvron et al. (2023)
22 CTRL Keskar et al. (2019) Text Wikipedia, Project Gutenberg, OpenWebText Gender, racial biases Gehman et al. (2020)
23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020); Nakamura et al. (2024)
24 Mixtral-8x7B Jiang et al. (2024) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024); Aich et al. (2024)
25 BLOOM Le Scao et al. (2023) Text ROOTS Toxicity, gender, religious, disability, age biases Le Scao et al. (2023)
The introduction of the Generative Pretrained Transformer (GPT-2) Radford et al. (2019) was a turning point in the LLM lanscape with its 1.5B parameters. As pointed out earlier, training such a large model required a lot of data and the Internet-sourced WebText was used for this purpose. Updated versions of the dataset have also been used for its recent successors, including GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023). The attendant problems of bias witnessed in GPT-2 Gehman et al. (2020) have followed successive versions.
Quantitative bias measurement in NLP may be placed into 3 categories: preuse, as carried out by Alkhaled et al. (2023) and Adewumi et al. (2023b), intrinsic Caliskan et al. (2017); May et al. (2019), and extrinsic, as observed by Ramesh et al. (2023). The first, second and third involve quantifying bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively. Far more work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al. (2023). For example, Delobelle et al. (2022) and Welbl et al. (2021) measured bias in pre-trained language models. Others have surveyed fairness and bias in LLM s Bender et al. (2021); Meade et al. (2022); Gallegos et al. (2023); Chang et al. (2024); Myers et al. (2024).
In other works, Schramowski et al. (2022) show that moral directions, i.e. what is morally right or wrong to do, are present in LLM s. It is, however, highly debatable if models can be considered moral agents. Also, Santurkar et al. (2023) explored whose opinions LLM s reflect. With regards to healthcare and bias, work has also been done Harrer (2023); Nashwan and Abujaber (2023). In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias.
Bias in multilingual AI:
Some of the multimodal data in Table 2 contain multilingual data. Such result in multilingual models and embeddings. For example, CC3M-35L, COCO-35L, and WebLI, used to train PaLI. WebLI is a mix of pre-training tasks with texts in 109 languages Chen et al. (2022). Some of these languages fall in the category of low-resource languages Adewumi et al. (2023a). Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words, using WEAT Caliskan et al. (2017). In the study by Wambsganss et al. (2022), they found that the pretrained German language models, GermanBERT, GermanT5, and German GPT-2, had substantial conceptual, racial, and gender bias. This was also confirmed by Kraft et al. (2022), who observed sexist stereotypes in some of the models (e.g. family- and care-related terms related to female while crime and perpetrators related to male). Similarly, for Dutch, Delobelle et al. (2020) investigated gender and occupation biases in RobBERT (a Dutch RoBERTa Liu et al. (2019)), through a template-based association test Kurita et al. (2019); May et al. (2019). Huang et al. (2020) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
5 Discussion
Although there are many limitations or risks of multimodal AI or LLM s Acosta et al. (2022); Adewumi et al. (2024b); Pettersson et al. (2024); Adewumi et al. (2024a), perhaps, the issue of fairness and bias rank among the topmost Mehrabi et al. (2021). In addition, in the taxonomy of 21 risks of language models provided by Weidinger et al. (2022), the first category is "Discrimination, Hate speech and Exclusion". Given, the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al. (2018); Rudinger et al. (2018). It may be almost impossible to automatically filter a dataset or debias a model to be 100% free of unfair, bias or toxic content but the research community and other stakeholders may need to determine what levels are acceptable and if it should be a requirement to have human-in-the-loop methods. In this section, we discuss further methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies. We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free, as much as possible, from bias.
5.1 Methods to audit, measure, and evaluate fairness and bias
Embedding Association Tests (EATs)have been used in several studies Caliskan et al. (2017); Kurpicz-Briki (2020); Wolfe et al. (2023), starting with the introduction of Word Embedding Association Test (WEAT)by Caliskan et al. (2017) and improvements in Sentence Embedding Association Test (SEAT)May et al. (2019) and Relational Inner Product Association (RIPA)Ethayarajh et al. (2019). Despite its widespread use, WEAT has the disadvantage that it may systematically overestimate the bias in a model. Another embedding evaluation method is cosine similarity. It was used for zero-shot classification by Radford et al. (2021). It may also be used to audit fairness and bias by evaluating the similarity in image and text embeddings Wolfe et al. (2023). The visual tool, Gradient-weighted Class Activation Mapping (GRAD-CAM), generates a saliency map, which shows the most relevant regions of an image for given attributes Selvaraju et al. (2017); Wolfe et al. (2023). In an evaluation carried out by Wolfe et al. (2023), they discovered that the computed average saliency maps included only face regions for non-objectified images but both face and chest regions for objectified images, in a possible indication of sexual objectification bias. Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al. (2021); Birhane et al. (2024b)
A recent metric introduced by Alkhaled et al. (2023) is bipol. It uses a two-step procedure in estimating bias in data Adewumi et al. (2023b); Pagliai et al. (2024). Bipol has the weakness that if the bias classifier is not accurate enough, false positives will weaken the evaluation score. Another measure is Area Under the Curve (AUC), as used by Meng et al. (2022) in the investigation of algorithmic fairness of mortality prediction, where they noted that Machine Learning (ML)methods obtain lower scores, usually, when it involves groups with higher mortality rates. Teo et al. (2024) proposed CLassifier Error-Aware Measurement (CLEAM), a framework for better performance in bias estimation, while Booth et al. (2021) measured gender bias using accuracy of Spearman rank-based correlation ($\rho$). Furthermore, Nozza et al. (2021) introduced the score “HONEST”, which was tested with respect to gender bias in text generation in 6 languages: Italian, French, Portuguese, Romanian, Spanish and English. It measures the probability that a language model will output hurtful text given a certain template and lexicon.
5.2 Datasets for Bias Evaluation
Different datasets have been introduced for bias evaluation. Esiobu et al. (2023) introduced 2 novel datasets AdvPromptSet and HolisticBiasR, with which they evaluated 12 demographic dimensions for different LLM s. Ruzzante et al. (2022) introduced Sexual OBjectification and EMotion Database (SOBEM)for sexual objectification bias studies. It consists of 280 pictures of objectified and non-objectified female models with 3 different emotions and a neutral face. Bias Benchmark for QA (BBQ)is a question-sets dataset that employs templates crafted to reflect specific biases identified in society. It was introduced by Parrish et al. (2022) and aims to expose implicit prejudices that may exist against individuals from legally protected categories. BEAVERTAILS was introduced by Ji et al. (2024). It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exclusive. Discrimination, Stereotype, Injustice make up the second category in the list. Additional datasets for bias evaluation include RedditBias by Barikeri et al. (2021), RealToxicityPrompts, which comprises of 100K English sentences Gehman et al. (2020), HarmfulQ for zero-shot Chain of Thought (CoT)across stereotype benchmarks and harmful questions Shaikh et al. (2023), and Bold Dhamala et al. (2021).
5.3 Debiasing strategies on data, algorithms, procedures
5.3.1 Curate over Crawl
One important method to address bias in datasets will be to curate over crawl. This is especially so because web crawling has been the popular approach to getting internet data in the shortest possible time Birhane et al. (2021). The assumption of scale beats noise is the rationale for this approach by some researchers Jia et al. (2021). Unfortunately, this misconception about scaling does not only scale the quality part of the dataset but the noise along with it, no matter how small, as shown by Birhane et al. (2024b) when they observed 12% increase in hate content due to scaling. Clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
5.3.2 Counterfactual Data Augmentation (CDA)
Zhao et al. (2018) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods. It involves generating alternate examples of what exists (counterfactual) of data points to counter or mitigate bias. This method has gained attention in the field Meade et al. (2022); Barikeri et al. (2021).
5.3.3 Improved Filtering
It has been shown that poor filtering during the data creation process allows low quality data in the final dataset Birhane et al. (2024a). It may not be possible to automatically filter large data to be 100% fit for purpose but improving the existing methods of filtering will go a long way in mitigating bias.
5.3.4 Other strategies
Oversampling under-represented groups in the data has been advocated by Buolamwini and Gebru (2018). There are generation detoxifying methods with the potential to reduce bias Gehman et al. (2020). These include data-based further pretraining and decoding-based generation. Similarly, Cabello et al. (2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks. In another work, Yan et al. (2020) used adversarial learning. They added a discriminator to jointly learn with the predictor for the sensitive attributes. Meanwhile, Zayed et al. (2024) addressed fairness by pruning in LLM s while Guo et al. (2022) introduced auto-bias, by directly probing the biases in pretrained models through prompts. Liang et al. (2021) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s. Additionally, there are Self-Debias Schick et al. (2021), Hard-Debias Bolukbasi et al. (2016), SentenceDebias Liang et al. (2020) and Dropout methods Webster et al. (2020); Meade et al. (2022)
6 Conclusion
Fairness and bias are very important considerations in multimodal AI. In this work, we presented the challenges of fairness and bias in multimodal data, LMM s, and LLM s, defining what both terms mean within the scope of this survey, while acknowledging other definitions in the literature. We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across 17 publishers, which reveals the gap of study in large multimodal AI compared to LLM s, which this work contributes to filling. Our discussion around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues. For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating fairness and bias concerns, respectively.
Acknowledgements
This work is funded by the EU HumaneAI Project. The work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU).

### v2

@@
 The importance of addressing fairness and bias in artificial intelligence (AI)systems cannot be over-emphasized.
-Mainstream media has been awashed with news of incidents around stereotypes and bias in many of these systems in recent years.
-In this survey, we fill a gap with regards to the minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models along with the challenges affecting them; we identify a new category of quantifying bias (preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic; we critically discuss the various ways researchers are addressing these challenges.
-Our method involved two slightly different search queries on Google Scholar, which revealed that 33,400 and 538,000 links are the results for the terms "Fairness and bias in Large Multimodal Models" and "Fairness and bias in Large Language Models", respectively.
-We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenge of fairness and bias in multimodal AI.
+Mainstream media has been awashed with news of incidents around stereotypes and other types of bias in many of these systems in recent years.
+In this survey, we fill a gap with regards to the relatively minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models related to both types of AI along with the challenges of bias affecting them.
+We discuss the less-mentioned category of mitigating bias, preprocessing (with particular attention on the first part of it, which we call preuse).
+The method is less-mentioned compared to the two well-known ones in the literature: intrinsic and extrinsic mitigation methods.
+We critically discuss the various ways researchers are addressing these challenges.
+Our method involved two slightly different search queries on two reputable search engines, Google Scholar and Web of Science (WoS), which revealed that for the queries ‘Fairness and bias in Large Multimodal Models’ and ‘Fairness and bias in Large Language Models’, 33,400 and 538,000 links are the initial results, respectively, for Scholar while 4 and 50 links are the initial results, respectively, for WoS.
+For reproducibility and verification, we provide links to the search results and the citations to all the final reviewed papers.
+We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenges of fairness and bias in multimodal and language AI.
 1 Introduction
@@
 (2024).
-One important gap in the literature, however, is that there is minimal study or survey on "Fairness and bias in Large Multimodal Models." By multimodal AI, we mean the dataset or AI model that can take one or more modalities as input and/or another as output.
-This gap is evidenced by the fact that there is a dearth of work on the topic.
-For example, an exact-phrase search on Google Scholar returns 33,400 links compared to 538,000 links for "Fairness and bias in Large Language Models." [^1] June 10, 2024 This implies more than 16 times the result for the former.
-Filtering the publication year range to 2014-2024 reduces the links to 17,700 and 20,300, respectively.
+One important gap in the literature, however, is that there is relatively minimal study or survey on ‘Fairness and bias in Large Multimodal Models.’ By multimodal AI, we mean the datasets or AI models that can take one or more modalities as input and/or another as output.
+This gap is evidenced by the fact that there are fewer works around the topic.
+For example, a query search on Google Scholar returns 33,400 links compared to 538,000 links for ‘Fairness and bias in Large Language Models’ (where the first query search is equivalent to the boolean operation fairness AND and AND bias AND in AND large AND multimodal AND models).
+[^1] July 10, 2024 This implies more than 16 times the result compared to the former.
+However, filtering the publication year range to 2014-2024 reduces the links to 17,200 and 19,300, respectively.
 We intend to contribute in filling that gap in this work.
@@
 (2024).
-Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one group compared to another, given the ground truth Booth et al.
+Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one entity compared to another, given the ground truth Booth et al.
 (2021); Scheuneman (1979).
@@
 Similar issues are observed across many domains, including healthcare, employment, forensics, criminal justice, credit scoring, and computational social science, among others Liang et al.
-(2021); Ferrara (2023); Landers and Behrend (2023); Han (2023).
+(2021b); Ferrara (2023a); Landers and Behrend (2023); Han (2023).
 According to Wolfe et al.
@@
 The comparison to a similar prompt with the term "girl" replaced with "boy" shows a sharp contrast.
+In related work, many of the relatively recent surveys on fairness and bias in AI appear to have been of a general nature or focused on other areas.
+Pagano et al.
+(2023) focused their attention on ML generally and the 5-year period between 2017 and 2022, thereby missing the nuances and some of the details related to natural language processing (NLP)and multimodal AI.
+Mehrabi et al.
+(2021) surveyed applications that have exhibited bias in different domains, listed sources of bias in these applications and created a taxonomy for fairness definitions.
+On the other hand, Le Quy et al.
+(2022) paid attention to benchmark tabular datasets for fairness, analysing relationships between different protected and class attributes.
+Balayn et al.
+(2021) focused on data bias in data engineering and management research, arguing for the enforcement of fairness requirements and constraints on the data for training and evaluating systems.
+Their survey method is limited in that it restricted part of its literature search between 2019 and 2020.
+In our work, in addition to discussing datasets that make AI models biased and the datasets for evaluating bias and fairness, we discuss other important related areas, such as the mitigation strategies.
+Blodgett et al.
+(2020) surveyed over 140 articles about bias in NLP and realized that the stated motivations are usually inconsistent and vague, and the articles do not engage with the broader applicable literature that is external to NLP.
+They, therefore, made 3 recommendations for those in this field: (1) establish the relationships in language and social hierarchies by referring to the broader literature outside of NLP, (2) be explicitly clear on why the system described as biased is harmful, and (3) evaluate the language of people affected by the biased systems.
+Their survey was restricted to only articles on text-based NLP, thereby excluding speech or multimodal AI, and limited to articles before May 2020.
+However, we follow their recommendations and our work answers some of the research questions identified in their work.
+For example, Section 4 and subsection 5.2.1 address the question How are datasets collected?.
+The survey by Sun et al.
+(2019) focused on recognizing and mitigating NLP gender bias and the authors discussed this based on four forms of representation bias: denigration, stereotyping, recognition, and under-representation.
+In Haltaufderheide and Ranisch (2024), they identified ethical issues around fairness and bias with LLM s in medicine and healthcare.
+Additional works that have surveyed fairness and bias in LLM s include Bender et al.
+(2021); Meade et al.
+(2022); Gallegos et al.
+(2023); Chang et al.
+(2024); Myers et al.
+(2024).
 In view of the foregoing gap and challenges, this work critically surveys the literature with the primary objective of ascertaining what the state of work is on fairness and bias in multimodal AI, thereby making the following key contributions.
   1. 1.
-    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges affecting them.
+    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges of bias affecting them.
   2. 2.
-    We identify a new category of quantifying bias (i.e. preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic Ramesh et al. (2023); Cabello et al. (2023).
+    We discuss the less-mentioned category of mitigating bias in NLP (i.e.
+preprocessing), though more common in general ML Mehrabi et al.
+(2021); Pagano et al.
+(2023), with particular attention on the first part of it, which we call preuse.
+The other two well-known mitigation methods in the literature are the intrinsic (or in-processing) and extrinsic (or post-processing) methods Ramesh et al.
+(2023); Cabello et al.
+(2023).
   3. 3.
@@
 The rest of this paper is organised as follows.
-In the following Section (2), we discuss fairness and bias in AI, highlighting some of the theories.
+In the following Section (2), we highlight some of the theories around fairness and bias in AI.
 In Section 3, we explain the method used for this survey.
@@
 2 Fairness and Bias in AI
+2.1 Implicit and Explicit Bias
+There is distinction between implicit and explicit bias such that, given two sets of terms that express the bias axis (e.g.
+gender or race), one set of male gender terms could be $\{S_{1}\}$ = {dad, man} and the other set of female terms be $\{S_{2}\}$ = {mum, woman}.
+Implicit bias is sufficiently specified with both lists.
+However, explicit bias requires two additional sets of attributes $\{A_{1}\}$ = {engineer, doctor} and $\{A_{2}\}$ = {caregiver, carer} that express the terms to which the earlier gendered terms exhibit association, albeit to different levels.
+Hence, a gender biased system could result in male terms in $\{S_{1}\}$ being strongly associated with attributes of career terms $\{A_{1}\}$ compared to female terms $\{S_{2}\}$, which could be strongly associated with attributes of home-related terms $\{A_{2}\}$ Friedrich et al.
+(2021).
+Besides these two broad distinction of bias, there are many types of bias, depending on the philosophical or social perspectives that may be taken.
+Hence, a discussion about every possible type of fairness or bias is beyond the scope of this study but we refer readers to Mehrabi et al.
+(2021), Van der Wal et al.
+(2024) and Navigli et al.
+(2023) for some of the many types that may be listed.
 2.2 Concepts of Fairness and Bias
-A discussion about every possible type of fairness or bias is beyond the scope of this study, however, we refer readers to Mehrabi et al.
-(2021) for some of the discussion.
-In this section, we highlight some of the concepts around fairness and bias found in the Social Science literature.
+In this section, we highlight a few of the concepts around fairness and bias found in the Social Science literature.
 2.2.1 Justice Theory
@@
 (2023).
-2.2 Consequences & Legal Perspectives
+2.3 Consequences of Bias
 Fredrickson and Roberts (1997) confirms that objectification victimises the subject and may result in habitual body monitoring, thereby increasing mental health risks, sexual dysfunction, eating disorders, and depression.
@@
 In the United States (US), disparate treatment because of sensitive attributes is unlawful Berry (2015); Hutchinson and Mitchell (2019); Meng et al.
-(2022).
+(2022a).
 This is also the case in some other countries Zafar et al.
 (2017).
-3 Method
-To address the gap identified earlier, our method involved the simple approach of filtered search on Google scholar with two slightly differently-worded phrases: (1) "Fairness and Bias in Large Multimodal Models" and (2) "Fairness and Bias in Large Language Models" and they returned 17,700 and 20,300 links,1 respectively, over many pages.
-The search was filtered to the period 2014-2024 during which deep learning made significant progress.
-Apparently, there are fewer scientific papers on the former.
-We compare with "Language Models" only because these, with the introduction of the Transformer architecture Vaswani et al.
+3 Methodology
+In order to have a fair and thorough survey for the stated objective in Section 1, we followed the general guidelines recommended for conducting a systematic literature review, which is founded on a rigorous and auditable methodology Kitchenham (2004); Brereton et al.
+(2007).
+We used two common scientific search engines: Google Scholar and Web of Science (WoS).
+Both are advantageous because they index the main literature databases or publishers, including the Institute of Electrical and Electronics Engineers (IEEE), Association for Computing Machinery (ACM), Conference on Neural Information Processing Systems (NeurIPS), Multidisciplinary Digital Publishing Institute (MDPI), Public Knowledge Project (PKP), Massachusetts Institute of Technology Press (MIT Press), Public Library of Science (PLoS), Association of Computational Linguistics (ACL), AI Access Foundation (AIAF), Higher Education Press (HE Press), Association of Measurement and Evaluation in Education and Psychology (EPODDER), New York University School of Law (NYUSL), Academic Conferences and Publishing International (ACPI), Machine Learning Research Press (MLRP), Journal of Machine Learning Research (JMLR), and Society of Photographic Instrumentation Engineers (SPIE).
+WoS provides a second assessment to Scholar because of the limitations of the latter, including its inclusion of non-peer-reviewed links or possible predatory journals in its results of primary studies (or papers) Foltỳnek et al.
+(2019).
+The two search engines are identified as largely reliable, helping to mitigate the risk of incompleteness in the search Foltỳnek et al.
+(2019).
+For both, we employed a similar multi-phase search process itemized below, which helped to bring refinement to each search and the final results.
+We are confident this approach returned the most applicable results for the purpose of this study.
+  1. 1.
+    In the first phase, we designed the following two search queries based on our objective.
+    • Fairness and Bias in Large Multimodal Models
+    • Fairness and Bias in Large Language Models
+    The boolean operation, in both search engines, for the first is equivalent to fairness AND and AND bias AND in AND large AND multimodal AND models while the second is fairness AND and AND bias AND in AND large AND language AND models.
+  2. 2.
+    In the second phase, we filtered the results based on the time period of just over 10 years (2014 to 2024) and relevance (after a review of paper titles and abstracts for over 200 papers).
+For example, one result involved an article about ‘executive coaching programs’ that has nothing to do with AI or social bias.
+All the filtered papers are in English.
+In addition, we corrected for misplaced results (i.e.
+papers that turned up in one search result though they belong in the other; about 15 in multimodal belonging in language and 9 in language belonging in multimodal, for Scholar).
+We removed papers published in journals in Beall’s List of Potential Predatory Journals and Publishers,222beallslist.net/standalone-journals if present.
+  3. 3.
+    In the third phase, we critically reviewed the papers for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
+More concretely, for WoS, we searched ‘All Fields’ of the documents ‘Core Collection’ across all ‘Editions’ without initially restricting the year of publication.
+For Scholar, the first phase returned 33,400 and 538,000 result links over many pages for the first and second queries, respectively, while returning 4 and 50 for WoS.1 Filtering Scholar, based on the time period, returned 17,200 and 19,300 links, which finally reduced to 69 and 101 at the end of the second phase for the first and second terms, respectively, while returning 8 and 44 for WoS.
+It is noteworthy that for Scholar, there were equally 100 links to start with for both queries.
+The boolean search operation involved all the individual words in any arbitrary order without case sensitivity and narrowed the search to mostly relevant documents.
+We compare ‘Multimodal Models’ with only ‘Language Models’ in the search terms because the latter, with the introduction of the Transformer architecture Vaswani et al.
 (2017), have influenced computer vision Yuan et al.
-(2021) and serve as core component for many multimodal AI.
-We captured archived papers (e.g.
-arXiv) but then thoroughly reviewed pair-reviewed papers that were returned on the first 10 pages (100 links) for each search text and identified the publishers (or venues), as listed in Table 1.
-We note that the more relevant papers turn up on the first page of the search while the quality of results degrade as one progresses through the pages.
-From the Table, it can be observed that there are 36 pair-reviewed papers on multimodal compared to 39 on language, among the 200 papers.
-We note that Google Scholar sometimes returns the arXiv version of a paper instead of the peer-reviewed version, hence, it’s very useful to include the data from arXiv.
-In total, 55 papers were published under multimodal compared to 88 for language.
-Each paper was reviewed for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
-Table 1: Distribution of scientific papers (2014-2024)
+(2021) and they serve as important components for many multimodal AI.
+Furthermore, we recognize that multimodal implies the combination of ‘language’ (or ‘speech’) and ‘vision’ but we did not use this distinction in our multimodal search because (1) the boolean equivalent can result in false positives, (2) even when we attempted it, we had less than 50% links in result compared to the second term of only ‘language’ search, and (3) ‘multimodal’ is a standard term in the field.
+Table 1: Distribution of scientific papers on Google Scholar. (Filtered total is limited to the year range 2014 - 2024 and to the first 100 links per search query, excluding irrelevant links.)
 # Publisher Multimodal Language
-1 IEEE 11 2
-2 ELSEVIER 2 2
-3 ACM 9 9
-4 Springer 1 1
-5 NeurIPS 4 8
-6 Nature 3 1
-7 MDPI - 1
-8 MLR 3 5
-9 PKP 2 -
+Unfiltered total 33,400 538,000
+1 IEEE 10 -
+2 Elsevier 3 3
+3 ACM 14 13
+4 Springer 7 5
+5 NeurIPS 4 11
+6 Nature 3 5
+7 MDPI 2 1
+8 MLRP 2 4
+9 PKP 1 1
 10 MIT Press - 4
-11 PubMed 1 -
+11 PubMed 1 1
 12 Cambridge - 1
-13 Patterns - 1
-14 PLOS - 1
+13 PLoS 1 1
 14 JMLR - 2
-16 Science - 1
-Sub-total 36 39
-17 arXiv 19 49
-Total 55 88
-4 Fairness and Bias in LLM s and LMM s
-In this section we review some LMM s and LLM s and the fairness and bias challenges they have.
-Tables 2 and 3 summarize some relevant datasets and the models, respectively.
-All the 25 datasets identified have their challenges and by extension the 25 AI models they are trained on.
+15 ACL 5 23
+16 SPIE 1 -
+17 De Gruyter 1 -
+18 Wiley 1 1
+19 Sage 1 -
+20 Academic Pinnacle 1 1
+Filtered sub-total 58 77
+21 arXiv 11 23
+22 Preprints - 1
+Filtered total 69 101
+Table 2: Distribution of scientific papers on Web of Science (WoS). (Filtered total is limited to the year range 2014 - 2024 and removes irrelevant results.)
+# Publisher Multimodal Language
+Unfiltered total 4 50
+Corrected total 8 46
+1 IEEE 4 5
+2 Elsevier - 2
+3 ACM 3 12
+4 Springer - 5
+5 NeurIPS - 1
+6 Nature - 1
+7 MDPI - 3
+22 MLRP - 1
+22 PKP 1 -
+16 ACL - 10
+13 Wiley - 2
+17 AIAF - 1
+12 Now - 1
+Filtered total 8 44
+Finally, for Scholar, we captured archived papers (e.g.
+arXiv) because, sometimes, their peer-reviewed versions exist but may not appear in the result.
+The summary statistics of the search are presented in Tables 1 and 2 while the references to the papers and the link to the search results are provided in the Appendix.
+The useful data from the reviewed papers (or primary studies) about datasets, AI models, and other contributions are then discussed in Sections 4.1, 4.2, and 5, respectively.
+We note that the more relevant papers turn up on the first page or at the top of the search while the quality of results degrade as one progresses through the pages or list.
+From the Tables, it can be observed that there are fewer pair-reviewed papers on multimodal models compared to language models.
+4 Findings on Fairness and Bias in LMM s and LLM s
+It is commonly agreed that AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022a).
+Tables 3 and 4 summarize some relevant datasets and the models beset by challenges of bias, respectively.
+All the 25 datasets identified have their challenges and by extension the 25 AI models which train on them.
 Some of these challenges include stereotypes, porn, misogyny, racial, gender, religious, cultural, age, and demographic biases.
-It is commonly agreed AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022).
 4.1 LMM s
-Table 2: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the second part of the table are usually used in downstream tasks).
-# Dataset Modality Some Challenges
+Table 3: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the lower part of the table are usually used in downstream tasks).
+# Dataset Modality Some Challenges of Bias/Fairness
 1 CommonCrawl Raffel et al. (2020) Text & Vision Fake news, hate speech, porn & racism Gehman et al. (2020); Luccioni and Viviano (2021)
@@
 9 Colossal Cleaned CommonCrawl (C4) Raffel et al. (2020) Text Offensive language, racial bias Raffel et al. (2020)
-10 The Pile Gao et al. (2020) Text Religious, racial, gender biases Gao et al. (2020)
+10 The Pile Gao et al. (2020a) Text Religious, racial, gender biases Gao et al. (2020a)
 11 CCAligned El-Kishky et al. (2020) Text Porn, racial bias El-Kishky et al. (2020)
@@
 21 MS COCO Lin et al. (2014) Text & Vision Gender bias Cabello et al. (2023); Zhao et al. (2017)
-22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022)
-23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022)
+22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022a)
+23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022a)
 24 MAB Alkhaled et al. (2023) Text Racism, misogyny, stereotypes Alkhaled et al. (2023); Pagliai et al. (2024)
-25 Twitter corpus Huang et al. (2020) Text Age, gender, racial biases Huang et al. (2020)
-Wolfe et al.
+25 Twitter corpus Huang et al. (2020b) Text Age, gender, racial biases Huang et al. (2020b)
+Liang et al.
+(2021a) acknowledged that multimodal representations are challenging because they seek to integrate information from multiple areas in applications like robotics, finance, healthcare and more.
+However, multimodal AI has not enjoyed enough resources to study generalization across different modalities and the complexities of training.
+With regards to bias, Wolfe et al.
 (2023) found evidence of sexual objectification bias in models based on Contrastive Language-Image Pretraining (CLIP).
@@
 (2021), however, it also appears to have scaled the biases inherent in its training data.
-Also, Wolfe and Caliskan (2022) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al.
+Also, Wolfe and Caliskan (2022a) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al.
 (2021), SLIP Mu et al.
 (2022), and BLIP Li et al.
-(2022), as measured with EATs.
-For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used Chicago Face Database (CFD)Ma et al.
+(2022), as measured with Embedding Association Tests (EATs).
+For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used the Chicago Face Database (CFD)Ma et al.
 (2015).
@@
 Similarly, Teo et al.
-(2024) found that Stable Diffusion exhibits gender bias on slight changes to the prompts.
-Besides work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022).
+(2024) found that Stable Diffusion exhibits gender bias on slight changes to its prompts.
+Besides the work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022).
 In the work by Fenu and Marras (2022), they perform comparative analysis on audio-visual speaker recognition systems, using fusion at the model step.
@@
 (2021) performed a case study of automated video interviews and found that combining more than one modality increases bias and reduces fairness, similarly to what happens when scaling crawled data for training models.
-Cabello et al.
-(2023) investigated fairness and bias in LMM s.
 Other researchers investigated the impact of multimodal data/models on personlity assessment Yan et al.
@@
 (2020), health records Meng et al.
-(2022) and more Birhane et al.
-(2021); Ferrara (2023); Zhang et al.
-(2022)
-4.2 LLM s
+(2022a) and more Birhane et al.
+(2021); Zhang et al.
+(2022c); Ferrara (2023a); Cabello et al.
+(2023) The obvious challenges of bias in data has motivated some researchers for more attention in the ethics of data collection Weinberg (2022).
 Table 4: Summary of Models and Some of Their Fairness & Bias Challenges. (*modified for audio-visual)
-# LMM s Modality Training Data Some Challenges
-1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022)
-2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023)
+# LMM s Modality Training Data Some Challenges of Bias/Fairness
+1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022a)
+2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023a)
 3 GLIDE Nichol et al. (2022) Text & Vision WIT & Conceptual Captions Gender stereotypes Nichol et al. (2022)
 4 Stable Diffusion Rombach et al. (2022) Text & Vision LAION-5B Gender bias Teo et al. (2024)
-5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022)
-6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022)
-7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022); Ruggeri and Nozza (2023)
-8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language222ai.google.dev/gemma/docs/paligemma/model-card
+5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022a)
+6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022a)
+7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022a); Ruggeri and Nozza (2023)
+8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language333ai.google.dev/gemma/docs/paligemma/model-card
 9 PaLI-3 Chen et al. (2022) Text & Vision WebLI Age, racial, gender biases & stereotypes Chen et al. (2022)
 10 Falcon 2 Almazrouei et al. (2023) Text & Vision RefinedWeb Harmful content, cultural bias Almazrouei et al. (2023)
-11 BEiT Wang et al. (2023) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
+11 BEiT Wang et al. (2023b) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
 12 LLaVA Liu et al. (2024b, a) Text & Vision Conceptual Captions Cultural bias Liu et al. (2024b)
@@
 14 GPT4o Achiam et al. (2023) Text, Audio & Vision WebText, Github, etc Stereotypes, racial bias Aich et al. (2024)
-LLM s
 15 GPT3 Brown et al. (2020) Text CommonCrawl & WebText Gender, racial, religious biases Brown et al. (2020); Gehman et al. (2020)
@@
 22 CTRL Keskar et al. (2019) Text Wikipedia, Project Gutenberg, OpenWebText Gender, racial biases Gehman et al. (2020)
-23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020); Nakamura et al. (2024)
-24 Mixtral-8x7B Jiang et al. (2024) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024); Aich et al. (2024)
+23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020a); Nakamura et al. (2024)
+24 Mixtral-8x7B Jiang et al. (2024a) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024a); Aich et al. (2024)
 25 BLOOM Le Scao et al. (2023) Text ROOTS Toxicity, gender, religious, disability, age biases Le Scao et al. (2023)
+4.2 LLM s
+The relevant datasets, models, and challenges of bias affecting LLM s, as discussed in the literature, are summarized in Tables 3 and 4.
 The introduction of the Generative Pretrained Transformer (GPT-2) Radford et al.
-(2019) was a turning point in the LLM lanscape with its 1.5B parameters.
+(2019) was a turning point in the language model landscape with its 1.5B parameters.
 As pointed out earlier, training such a large model required a lot of data and the Internet-sourced WebText was used for this purpose.
@@
 (2023).
-The attendant problems of bias witnessed in GPT-2 Gehman et al.
+Unfortunately, the attendant problems of bias witnessed in GPT-2 Gehman et al.
 (2020) have followed successive versions.
-Quantitative bias measurement in NLP may be placed into 3 categories: preuse, as carried out by Alkhaled et al.
-(2023) and Adewumi et al.
-(2023b), intrinsic Caliskan et al.
+This is also the case with other models, as expressed in Table 4, for the reasons that the datasets used in pretraining, some of which are given in Table 3, are largely from Internet sources.
+Furthermore, the architectures of the models, many of which are based on the Transformer architecture Vaswani et al.
+(2017), share similarity.
+In other works, Schramowski et al.
+(2022) show that moral directions, i.e.
+what is morally right or wrong to do, are present in LLM s.
+It is, however, highly debatable if models can be considered moral agents.
+In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias.
+Also, Santurkar et al.
+(2023) explored whose opinions LLM s reflect while Harrer (2023); Nashwan and Abujaber (2023) investigated bias in LLM s within healthcare.
+4.3 Mitigation Categories
+Quantitative bias measurement and mitigation in NLP may be placed into 3 categories: preprocessing, as applicable in general ML Hort et al.
+(2024), intrinsic Caliskan et al.
 (2017); May et al.
@@
 (2023).
-The first, second and third involve quantifying bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively.
-Far more work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al.
+The first, second and third involve quantifying and mitigating bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively.
+More work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al.
 (2023).
@@
 (2021) measured bias in pre-trained language models.
-Others have surveyed fairness and bias in LLM s Bender et al.
-(2021); Meade et al.
-(2022); Gallegos et al.
-(2023); Chang et al.
-(2024); Myers et al.
+One main reason why more work has been on the latter, according to Sun et al.
+(2019), is that debiasing an existing model to adjust the outputs usually only requires ‘patching’ the model instead of retraining with modified (debiased) or new data, which is usually costly.
+Preuse
+This may be considered the first step of the NLP preprocessing method for debiasing.
+It involves quantifying the amount of bias (or toxicity) in a given dataset by using a model, after the model has been trained on a different dedicated dataset, without necessarily attempting to mitigate the bias in the given dataset.
+It can be defined by Equation 1, where $B(d)$ is a bias metric that takes data as input and returns a scalar, $s$, as the score.
+Examples of works involving this are Alkhaled et al.
+(2023); Adewumi et al.
+(2023b).
+HateBERT Caselli et al.
+(2021) has been used in similar settings.
+Equation 1 is similar to that in Brunet et al.
+(2019) of differential bias, where they approximate the effect of removing small parts of training data on bias.
+Gender bias is one example of the kind of bias that can be estimated, as it has been observed that there are more male than female terms in many NLP datasets Sun et al.
+(2019); Alkhaled et al.
+(2023); Pagliai et al.
 (2024).
-In other works, Schramowski et al.
-(2022) show that moral directions, i.e.
-what is morally right or wrong to do, are present in LLM s.
-It is, however, highly debatable if models can be considered moral agents.
-Also, Santurkar et al.
-(2023) explored whose opinions LLM s reflect.
-With regards to healthcare and bias, work has also been done Harrer (2023); Nashwan and Abujaber (2023).
-In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias.
-Bias in multilingual AI:
+The ability to first estimate quantitatively the bias in a given dataset provides the basis to be able to determine the level of success of mitigation.
+$\mathit{B{\left(d\right)}}={s}$ (1)
+4.4 Bias in multilingual AI:
 Some of the multimodal data in Table 3 contain multilingual data.
-Such result in multilingual models and embeddings.
-For example, CC3M-35L, COCO-35L, and WebLI, used to train PaLI.
+These result in multilingual models and embeddings.
+For example, CC3M-35L, COCO-35L, and WebLI, which was used to train PaLI.
 WebLI is a mix of pre-training tasks with texts in 109 languages Chen et al.
@@
 (2023a).
-Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words, using WEAT Caliskan et al.
+Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words using WEAT Caliskan et al.
 (2017).
@@
 (2022), who observed sexist stereotypes in some of the models (e.g.
-family- and care-related terms related to female while crime and perpetrators related to male).
+family- and care-related terms were associated with female while crime and perpetrators were associated with male).
 Similarly, for Dutch, Delobelle et al.
@@
 Huang et al.
-(2020) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
+(2020b) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
 5 Discussion
@@
 (2024); Adewumi et al.
-(2024a), perhaps, the issue of fairness and bias rank among the topmost Mehrabi et al.
+(2024a), perhaps the issue of fairness and bias rank among the topmost Mehrabi et al.
 (2021).
 In addition, in the taxonomy of 21 risks of language models provided by Weidinger et al.
-(2022), the first category is "Discrimination, Hate speech and Exclusion".
-Given, the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al.
-(2018); Rudinger et al.
+(2022), the first category is ‘Discrimination, Hate speech and Exclusion’.
+Given the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al.
+(2018a); Rudinger et al.
 (2018).
 It may be almost impossible to automatically filter a dataset or debias a model to be 100% free of unfair, bias or toxic content but the research community and other stakeholders may need to determine what levels are acceptable and if it should be a requirement to have human-in-the-loop methods.
-In this section, we discuss further methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies.
-We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free, as much as possible, from bias.
+In this section, we discuss methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies.
+We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free from bias, as much as possible.
 5.1 Methods to audit, measure, and evaluate fairness and bias
-Embedding Association Tests (EATs)have been used in several studies Caliskan et al.
-(2017); Kurpicz-Briki (2020); Wolfe et al.
-(2023), starting with the introduction of Word Embedding Association Test (WEAT)by Caliskan et al.
-(2017) and improvements in Sentence Embedding Association Test (SEAT)May et al.
+Caliskan et al.
+(2017) introduced the Word Embedding Association Test (WEAT), which is based on the Implicit Association Test (IAT)Nosek et al.
+(2002).
+The IAT was designed to measure attitudes towards social groups.
+It showed implicit preference for White and young people over Black and old people, respectively.
+Furthermore, it showed the association of male terms with science while female terms were with family and arts.
+Embedding Association Tests (EATs)have been used in several studies Kurpicz-Briki (2020); Wolfe et al.
+(2023) and adapted with improvements in Sentence Embedding Association Test (SEAT)May et al.
 (2019) and Relational Inner Product Association (RIPA)Ethayarajh et al.
@@
 Despite its widespread use, WEAT has the disadvantage that it may systematically overestimate the bias in a model.
+In vision models, Mandal et al.
+(2023b) used WEAT to audit CLIP by detecting and quantifying bias.
+Also, along the lines of the WEAT, Dev and Phillips (2019) introduced the Embedding Coherence Test (ECT)and Embedding Quality Test (EQT)and proposed methods for eliminating explicit bias.
+However, their method has the weakness that it is not able to remove implicit bias Friedrich et al.
+(2021).
 Another embedding evaluation method is cosine similarity.
@@
 (2023), they discovered that the computed average saliency maps included only face regions for non-objectified images but both face and chest regions for objectified images, in a possible indication of sexual objectification bias.
-Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al.
+The tool Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al.
 (2021); Birhane et al.
@@
 Another measure is Area Under the Curve (AUC), as used by Meng et al.
-(2022) in the investigation of algorithmic fairness of mortality prediction, where they noted that Machine Learning (ML)methods obtain lower scores, usually, when it involves groups with higher mortality rates.
+(2022a) in the investigation of algorithmic fairness of mortality prediction, where they noted that ML methods obtain lower scores, usually, when it involves groups with higher mortality rates.
 Teo et al.
@@
 It measures the probability that a language model will output hurtful text given a certain template and lexicon.
-5.2 Datasets for Bias Evaluation
+Datasets for Bias Evaluation
 Different datasets have been introduced for bias evaluation.
+Liang et al.
+(2021a) introduced MultiBench, a unified multimodal benchmark that spans 15 datasets, 10 modalities, and 20 prediction tasks.
+FairFace was introduced by Karkkainen and Joo (2021).
+The dataset was designed to mitigate racial bias in multimodal AI, collected from the Yahoo Flickr Creative Commons 100 Million (YFCC100M)dataset Thomee et al.
+(2016) and contains 108,501 images, balanced across the following races: Black, White, Indian, Southeast Asian, East Asian, Middle Eastern, and Latino.
 Esiobu et al.
@@
 It consists of 280 pictures of objectified and non-objectified female models with 3 different emotions and a neutral face.
-Bias Benchmark for QA (BBQ)is a question-sets dataset that employs templates crafted to reflect specific biases identified in society.
+Bias Benchmark for QA (BBQ)is a question-set dataset that employs templates crafted to reflect specific biases identified in society.
 It was introduced by Parrish et al.
@@
 (2024).
-It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exclusive.
-Discrimination, Stereotype, Injustice make up the second category in the list.
+It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exhaustive.
+‘Discrimination, Stereotype, Injustice’ makes up the second category in the list.
 Additional datasets for bias evaluation include RedditBias by Barikeri et al.
@@
 (2021).
-5.3 Debiasing strategies on data, algorithms, procedures
+Porgali et al.
+(2023) also introduced the Casual Conversations dataset (version 2) containing 26,467 videos of 5,567 unique participants from 7 different countries, representing a wide range of demographics, for bias and robustness evaluation of LMM s that are vision and audio models.
+5.2 Debiasing strategies
+Although there’s no silver bullet to solving the challenges of fairness and bias in the data and models of multimodal AI, we believe a combination of two or more of the following strategies on the relevant datasets or models in Tables 3 and 4 will go a long way in mitigating bias in AI generally.
 5.2.1 Curate over Crawl
-One important method to address bias in datasets will be to curate over crawl.
+One important method to address bias in datasets will be to curate rather than crawl.
 This is especially so because web crawling has been the popular approach to getting Internet data in the shortest possible time Birhane et al.
@@
 (2024b) when they observed 12% increase in hate content due to scaling.
-Clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
+On the other hand, clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
+In addition, for better quality data collection, researchers like Jo and Gebru (2020) advocate that AI practitioners should build on the practices of those in the field of archives and libraries and have a public mission statement to guide their data collection practice Weinberg (2022).
+Furthermore, Gebru et al.
+(2021) encourage, through ‘datasheets for datasets’, standardized processes for documenting important aspects of the dataset creation, including motivation, composition, funding, collection and use cases, with the potential to mitigate unwanted biases, though this approach has its limitations because individuals included or affected by the datasets are not necessarily empowered to influence them Weinberg (2022).
 5.2.2 Counterfactual Data Augmentation (CDA)
 Zhao et al.
-(2018) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods.
+(2018a) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods.
 It involves generating alternate examples of what exists (counterfactual) of data points to counter or mitigate bias.
@@
 (2022); Barikeri et al.
+(2021).
+It is sometimes called ‘gender-swapping’ in the specific case of gender bias Sun et al.
+(2019).
+CDA has been shown to be effective in tackling bias in coreference resolution, as mentioned earlier, as it reduced the difference in F1 scores between pro-stereotypical and antistereotypical evaluations Sun et al.
+(2019).
+Dev and Phillips (2019) also embraced this neutralizing approach in their work by flipping gendered terms with their counterparts.
+For example, a sentence like ‘he was a doctor’ would be flipped to ‘she was a doctor’.
+Despite the advantages, as pointed out earlier about the less-discussed preprocessing bias mitigation method, completely debiasing textual data can be complicated and intricate.
+Some of the challenges are that the size of the data increases significantly, thereby increasing model training time, and naive term-swapping can create more challenges of unrealistic or nonsensical samples in the data, e.g.
+‘she has hot flashes because of menopause’ to ‘he has hot flashes because of menopause’ Sun et al.
 (2019).
@@
 It may not be possible to automatically filter large data to be 100% fit for purpose but improving the existing methods of filtering will go a long way in mitigating bias.
+5.2.4 Linear Projection
+This method projects all words $w\in W$ orthogonally to the bias vector, ensuring the updated set has no component along the bias vector, $v_{B}$, as given in Equation 2 Bolukbasi et al.
+(2016); Dev and Phillips (2019).
+$\mathit{w\textsuperscript{´}}=\mathit{w-\pi_{B}{\left(w\right)}}=\mathit{w-\langle w,v_{B}\rangle v_{B}}$ (2)
+The span that results becomes less by 1 in the total dimensions, say from 300 to 299, which will have negligible effects on the generalizability of the embeddings.
+As an example with gender bias, subtraction with linear projection of gender terms from embeddings will make them close.
+Gendered word-pairs that have few word sense (e.g.
+he - she) and (him - her) can have close enough identifcal positions in the vector space after debiasing.
+5.2.5 Debiasing Word Embeddings
+Removing gender bias from word embeddings can take one of two approaches: (1) removing gender subspace Bolukbasi et al.
+(2016) and (2) learning gender-neutral embeddings Zhao et al.
+(2018b).
+The two approaches may not be adequate for embeddings that are not based on Euclidean space since cosine similarity will no longer apply Sun et al.
+(2019).
+The first approach modifies an embedding based on the combined properties of word embeddings that gender bias can be captured by a direction and neutral words are linearly separable from gendered words Bolukbasi et al.
+(2016) while the second approach preserves gender information in some dimensions but compels other dimensions of the word embedding to be free of such Zhao et al.
+(2018b).
+5.2.6 Adapters
+A post-processing method for bias mitigation that is based on Adapter modules Houlsby et al.
+(2019), called Debiasing with Adapter Modules (DAM), was introduced by Kumar et al.
+(2023).
+They encapsulate different bias mitigation functionalities and can be integrated when desired in a model, similarly to how AdapterFusion Pfeiffer et al.
+(2021) is carried out in multi-task learning.
+DAM trains the main adapter and the bias mititgation adapters independently before combining them.
+DAM follows an earlier adapter-based debiasing method, called Adapter-based DEbiasing of LanguagE Models (ADELE), performed in the work by Lauscher et al.
+(2021).
+Their approach involved the additional use of CDA.
+5.2.7 Additive Residuals
+To address the skewed distribution of different identity groups in the training data used in LMM s, Seth et al.
+(2023) introduced Debiasing with Additive Residuals (DeAR)to learn additive residual image representations.
+This minimizes the representations’ capacity to distinguish among different identity groups, thereby offering fairer output.
+5.2.8 Continued Pretraining
+Fatemi et al.
+(2023) built on the continued pretraining concept, which is sometimes used for gender bias mitigation with a small gender-neutral dataset de Vassimon Manela et al.
+(2021), by introducing GEnder Equality Prompt (GEEP)such that it reduces catastrophic forgetting, which is a likely event in continued pretraining Kirkpatrick et al.
+(2017).
+GEEP achieves this by freezing the entire model before updating the embeddings.
+Furthermore, Cabello et al.
+(2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks.
+5.2.9 Adversarial Learning
+Yan et al.
+(2020) used adversarial learning for bias mitigation as proposed by Zhang et al.
+(2018).
+They added a discriminator to jointly learn with the predictor for the sensitive attributes.
+In this approach, the generator prevents the discriminator from identifying gender in a task.
+5.2.10 Gender Tagging
+In Machine Translation (MT), gender-tagging may be used Vanmassenhove et al.
+(2018).
+It involves the addition of gender tags to the beginning of data samples to identify the gender of the source data, e.g.
+‘FEMALE I’m travelling tomorrow’.
+Apparently, for more complex sentences this approach may become more challenging Sun et al.
+(2019).
+5.2.11 FairDistillation
+To address bias across languages, Delobelle and Berendt (2022) introduced FairDistillation.
+It is a cross-lingual method that is based on knowledge distillation Hinton et al.
+(2015) by creating smaller language models from large ones to control for stereotypical and representational biases.
+5.2.12 DEBIE
+Friedrich et al.
+(2021) introduced DEBIE as a platform for measuring and mitigating implicit and explicit bias in word embeddings.
+The mitigation methods are more specific to NLP and not available in general purpose library such as AI Fairness 360 (AIF360)Bellamy et al.
+(2019).
+DEBIE is a collection of commonly used bias data tools and word embeddings, including fastText Bojanowski et al.
+(2017), GloVe Pennington et al.
+(2014), Continuous Bag-of-Words (CBoW)Mikolov et al.
+(2013), and the WEAT test.
 5.2.13 Other strategies
-Oversampling under-represented groups in the data has been advocated by Buolamwini and Gebru (2018).
-There are generation detoxifying methods with the potential to reduce bias Gehman et al.
+Furthermore, there are generation detoxifying methods with the potential to reduce bias Gehman et al.
 (2020).
-These include data-based further pretraining and decoding-based generation.
-Similarly, Cabello et al.
-(2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks.
-In another work, Yan et al.
-(2020) used adversarial learning.
-They added a discriminator to jointly learn with the predictor for the sensitive attributes.
-Meanwhile, Zayed et al.
+These include the earlier-mentioned continued pretraining and decoding-based generation.
+Buolamwini and Gebru (2018) advocated oversampling under-represented groups in data to mitigate bias.
+Zayed et al.
 (2024) addressed fairness by pruning in LLM s while Guo et al.
@@
 Liang et al.
-(2021) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s.
+(2021b) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s.
 Additionally, there are Self-Debias Schick et al.
@@
 (2020); Meade et al.
-(2022)
+(2022).
 6 Conclusion
@@
 In this work, we presented the challenges of fairness and bias in multimodal data, LMM s, and LLM s, defining what both terms mean within the scope of this survey, while acknowledging other definitions in the literature.
-We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across 17 publishers, which reveals the gap of study in large multimodal AI compared to LLM s, which this work contributes to filling.
-Our discussion around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues.
-For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating fairness and bias concerns, respectively.
+We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across many publishers, which reveals the gap in the study of large multimodal AI compared to LLM s, which this work contributes to filling.
+Our discussions around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues.
+For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating the challenges of fairness and bias because despite the positive effects of debiasing in AI, researchers like West et al.
+(2019) argue that such research ought to do more than technical debiasing and include the social analysis of the use of such AI, as this will account more for the impact of bias overall Weinberg (2022).
+Even Sami et al.
+(2023) showed with the example of DALLE-2 that current efforts still have their limitations.
+In their work on the LMM DALLE-2, they revealed that, despite the guardrails for the model by OpenAI, it generated 40 more images of women for a stereotypical female-dominant administrative task, in clear gender bias, when prompted.
 Acknowledgements
-This work is funded by the EU HumaneAI Project.
+This work is supported by the European Commission-funded project "Humane AI: Toward AI Systems That Augment and Empower Humans by Understanding Us, Our Society and the World Around Us." The authors wish to thank our colleagues at the Responsible Artificial Intelligence Group at Umeå University.
 The work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU).
+Appendix A Appendix
+A.1 Google Scholar Papers
+A.1.1 Multimodal444scholar.google.com/scholar?hl=en&as_sdt=0%2C5&q=Fairness+ and+Bias+in+Large+Multimodal+Models&btnG=
+IEEE
+: Wang et al.
+(2020), Niu et al.
+(2021a), Joshi et al.
+(2021), Wang et al.
+(2022c), Seth et al.
+(2023), Karkkainen and Joo (2021), Peng et al.
+(2022), Cho et al.
+(2023), Wang et al.
+(2022b), Niu et al.
+(2021b), Hirota et al.
+(2022b)
+Springer
+: Xu et al. (2020), Peña et al. (2023), Stoychev and Gunes (2022), Georgopoulos et al. (2021), Cai et al. (2024), Alwahaby et al. (2022), Kehrenberg et al. (2020)
+ACM
+: Booth et al.
+(2021), Yan et al.
+(2020), Wolfe and Caliskan (2022a), Goyal et al.
+(2022b), Cheong et al.
+(2024), Wolfe and Caliskan (2022b), Liang et al.
+(2022a), Yao et al.
+(2023), Hirota et al.
+(2022a), Yee et al.
+(2021), Rahman et al.
+(2020), Pessach and Shmueli (2022), Xu et al.
+(2024), Tang et al.
+(2021)
+arXiv
+: Luo et al.
+(2024), Zhou et al.
+(2021), Wang et al.
+(2024), Birhane et al.
+(2021), Friedrich et al.
+(2023), Goyal et al.
+(2022a), Zhang et al.
+(2024b), Rana and Jha (2022), Chen et al.
+(2024), Soares and Angelov (2019), Zong et al.
+(2022)
+PubMed
+: Liang et al. (2021a)
+MDPI
+: Ferrara (2023a), Nazi and Peng (2024)
+ACL
+: Wang et al. (2021), Srinivasan and Bisk (2022), Li et al. (2021), Zhang et al. (2024a), Ross et al. (2021)
+Nature
+: Meng et al. (2022a), Röösli et al. (2022), Fei et al. (2022)
+PLoS
+: Jiang et al. (2024b)
+SPIE
+: Drukker et al. (2023)
+De Gruyter
+: Jenks (2024)
+NeurIPS
+: Gadre et al. (2024), Liang et al. (2022c), Koh et al. (2024), Lee et al. (2023)
+Elsevier
+: Rahate et al. (2022), Serna et al. (2022), Kumari and Ekbal (2021)
+PKP
+: Jaiswal and Provost (2020)
+MLRP
+: Kiros et al. (2014), Zhang et al. (2022a)
+Wiley
+: Ntoutsi et al. (2020)
+Sage
+: Tay et al. (2022)
+A.1.2 Language666scholar.google.com/scholar?hl=en&as_sdt=0%2C5&q= Fairness+and+Bias+in+Large+Language+Models&btnG=
+NeurIPS
+: Ma et al.
+(2023), Vig et al.
+(2020), Kojima et al.
+(2022), Solaiman and Dennison (2021), Wick et al.
+(2019), Tan and Celis (2019), Ouyang et al.
+(2022), Schaeffer et al.
+(2024), Meng et al.
+(2022b), Brown et al.
+(2020), Yu et al.
+(2023)
+ACM
+: Bender et al.
+(2021), Chang et al.
+(2024), Dhamala et al.
+(2021), Weidinger et al.
+(2022), Mehrabi et al.
+(2021), Garg et al.
+(2019), Ganguli et al.
+(2022), Min et al.
+(2023), Guo and Caliskan (2021), Zhang et al.
+(2020), Dodge et al.
+(2019), Navigli et al.
+(2023), Kotek et al.
+(2023)
+PKP
+: Pryzant et al. (2020)
+MIT Press
+: Ziems et al. (2024), Jiang et al. (2021), Schick et al. (2021), Raza et al. (2024)
+Springer
+: Freiberger and Buchmann (2024), Meyer et al. (2023), Teubner et al. (2023), Hou et al. (2024), Lu et al. (2020)
+MLRP
+: Liang et al. (2021b), Biderman et al. (2023), Aher et al. (2023), Lehman et al. (2023), Liang et al. (2021b)
+JMLR
+: Chung et al. (2024), Chowdhery et al. (2023)
+Cambridge
+: Argyle et al. (2023)
+arXiv
+: Ferrara (2023b), Wang et al.
+(2023a), Tamkin et al.
+(2021), Chen et al.
+(2021), Liang et al.
+(2022b), Weidinger et al.
+(2021), bench authors (2023), Rae et al.
+(2021), Eloundou et al.
+(2023), Serapio-García et al.
+(2023), Qi et al.
+(2023), Smith et al.
+(2022b), Huang et al.
+(2023), Liu et al.
+(2023a), Liu et al.
+(2023b), Hu et al.
+(2021), Zhang et al.
+(2022b), Sinha et al.
+(2021), Thoppilan et al.
+(2022), Menick et al.
+(2022), Doan et al.
+(2024), weidinger2019ethical
+Elsevier
+: Kasneci et al. (2023), Liu et al. (2023c), Ray (2023)
+ACL
+: Welbl et al.
+(2021), Nadeem et al.
+(2021), Gehman et al.
+(2020), Delobelle et al.
+(2020), Sap et al.
+(2020), Park et al.
+(2018), Perez et al.
+(2022), Blodgett et al.
+(2020), Khanuja et al.
+(2023), Agrawal et al.
+(2022), Sun et al.
+(2019), Liang et al.
+(2020), Dinan et al.
+(2020), Blodgett et al.
+(2021), Chen and Mueller (2024), Gao et al.
+(2020b), Kurita et al.
+(2019), Yoo et al.
+(2021), Lin et al.
+(2022), Hutchinson et al.
+(2020), Ramesh et al.
+(2023), Smith et al.
+(2022a), Ruder et al.
+(2022)
+Nature
+: Schramowski et al. (2022), Clusmann et al. (2023), Hager et al. (2024), Meskó and Topol (2023), Thirunavukarasu et al. (2023)
+PLoS
+: Mozafari et al. (2020)
+Wiley
+: Hovy and Prabhumoye (2021)
+MDPI
+: Garrido-Muñoz et al. (2021)
+Preprints
+: Kumar et al. (2024)
+PubMed
+: Karabacak and Margetis (2023)
+Academic Pinnacle
+: Desai et al. (2023)
+A.2 Web of Science (WoS)Papers
+A.2.1 Multimodal [^8] https://www.webofscience.com/wos/woscc/summary/889a9d49-d906-408c-91b6-ffefdff1880d-fed7fb8c/relevance/1
+IEEE
+: Seth et al. (2023), Porgali et al. (2023), Sami et al. (2023), Petryk et al. (2022)
+ACM
+: Edenberg and Wood (2023), Mandal et al. (2023b), Alam (2022)
+PKP
+: Lui et al. (2024)
+A.2.2 Language101010www.webofscience.com/wos/woscc/summary/79cd811f-beb4-40f2-844f-e0ca9edf1fe2-fed7f83e/relevance/1
+IEEE
+: Malic et al. (2023), Li et al. (2024), Morales et al. (2023), Da et al. (2024), Jeong et al. (2024)
+Elsevier
+: Nazir et al. (2024), Fan et al. (2020)
+ACM
+: Dacon and Liu (2021), Ovalle et al.
+(2023), Halevy et al.
+(2021), Mei et al.
+(2023), Salinas et al.
+(2023), Dhamala et al.
+(2021), Gadiraju et al.
+(2023), Saxena et al.
+(2024), Jamil (2024), Zerveas et al.
+(2022), Wright et al.
+(2020), Giner-Miguelez et al.
+(2023), Raj et al.
+(2023)
+Springer
+: Dolci et al. (2023), Delobelle and Berendt (2022), Schröder et al. (2023), Leteno et al. (2023), Yang et al. (2022),
+NeurIPS
+: Ma et al. (2023)
+Nature
+: Haltaufderheide and Ranisch (2024)
+MDPI
+: Bevara et al. (2024), Piñeiro-Martín et al. (2023), da Silva et al. (2021)
+MLRP
+: Liang et al. (2021c)
+Now
+: Xue et al. (2023)
+Wiley
+: Lee et al. (2024), Hao et al. (2024)
+ACL
+: Kumar et al.
+(2023), Lauscher et al.
+(2021), Sheng et al.
+(2021), Fatemi et al.
+(2023), Guo et al.
+(2022), Vanmassenhove et al.
+(2018), Huang et al.
+(2020a), Felkner et al.
+(2023), Talat et al.
+(2022), Escudé Font and Costa-jussà (2019)
+AIAF
+: Van der Wal et al. (2024)